OpenStudy (anonymous):

Jeff is holding a deck of cards, and asks you to pick 2. What is the probability that you pick a heart the first time and a spade the second time? He says to put the first one back in the deck before picking the second one.

OpenStudy (unklerhaukus):

\[P(\color{red}\heartsuit , \spadesuit)=P(\color{red}\heartsuit )\times P(\spadesuit )\] \[P(\color{red}\heartsuit )=13/52=1/4\]\[P(\spadesuit )=13/52=1/4\]